Factor f(x) into linear factors given that k is a zero of f(x)
f(x)=3x³-10x²-37x+60; k=5
factor completely.

Answer by Edwin McCravy(20059)

Factor f(x) into linear factors given that k is a zero of f(x)
f(x)=3x³-10x²-37x+60; k=5
factor completely.

f(x)=3x³-10x²-37x+60


5|3  -10  -37  60
 |    15   25 -60 
  3    5  -12   0

f(x)=(x-5)(3x²+5x-12)

f(x)=(x-5)(3x-4)(x+3)

Edwin